The heights of male are normally distributed with mean of 170 cm and standard deviation of 7.5cm. Find the probability that a randomly selected male has a height > 180 cm.


1
Expert's answer
2022-04-06T15:38:59-0400

We have a normal distribution, "\mu=170, \sigma=7.5."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{180-170}{7.5}=1.33;"


"P(X>180)=P(Z>1. 33)=\\=1-P(Z<1.33)="

"=1-0.9082=0.0918" ​(from z-table).